Ten students are being interviewed for a class office. Six of them are female and four are male. Their names are all placed in a box and two students are selected for the interviews.
a. What is the probability that both of those selected are female?
b. What is the probability that at least one is male?

Answer by stanbon(57407)

Ten students are being interviewed for a class office. Six of them are female and four are male. Their names are all placed in a box and two students are selected for the interviews.
a. What is the probability that both of those selected are female?
Ans: 6C2/10C2 = 15/45 = 1/3
-------------------------------------
b. What is the probability that at least one is male?
P(at least one male) = 1 - P(both are female) = 1-1/3 = 2/3
===========================
Cheers,
Stan H.